Please help me solve this problem:
The solution of the system 4x+5y=2 and 6x-2y=b is (3,a). What are the values of a and b?

Answer by AnlytcPhil(1806)

Please help me solve this problem: 
The solution of the system 4x+5y=2 and 6x-2y=b is (3,a).
What are the values of a and b?

Since (3,a) is their common solution, it is a solution to both 
equations. So we substitute " 3 " for x and " a " for y in the 
first one:

    4x + 5y = 2
4(3) + 5(a) = 2
    12 + 5a = 2
         5a = -10
          a = -2

Now we substitute " 3 " for x and " a " for y in the second one:

    6x - 2y = b
6(3) - 2(a) = b
    18 - 2a = b

But since we have already found " a " to be -2, we substitute -2
for " a " and get

 18 - 2(-2) = b
     18 + 4 = b        
         22 = b

Edwin
AnlytcPhil@aol.com